OpenStudy (anonymous):

4=√-2 a.√6 b.36 c.3 d.6

OpenStudy (anonymous):

is there x to find or anything...??? square both sides...

OpenStudy (anonymous):

i thought u cant have negatives under a radical

OpenStudy (anonymous):

if there's no x or something ... they are not equal !

OpenStudy (anonymous):

it can't be solved, coz sq. root having a negative sign is undefined

OpenStudy (anonymous):

i meant 4=√p-2

OpenStudy (anonymous):

@tsitrond actually you can, is just that you have to do with complex numbers!! Example... \[\LARGE \sqrt{-25}=i\cdot 5\]

OpenStudy (anonymous):

then square both sides and you'll get 16=p-2 p=16+2 unless 2 is not under the square isn't it??

OpenStudy (anonymous):

\[4+2=\sqrt{p}\] 6=\[\sqrt{p}\]

OpenStudy (anonymous):

@kreshnik yes i know but that is unreal numbers... i assumed this was all real numbers situation